# Data-Driven Safety Verification using Barrier Certificates and Matrix Zonotopes

### v1

@@ V Numerical simulations
to incorporate the semi-algebraic set expressions for $w,A$ and $B$ for the optimization process based on condition (III). Figure 1 shows the resulting sublevel set of the barrier certificate for a two dimensional system along with the designated initial and unsafe states. Table I displays the true dynamics used for simulation (data collection) as well as the relevant vectors and matrices used for the SOS optimization.
Figure 1: Barrier Certificate for a 2D System.
Figure 2: Barrier Certificate Projections for a 5D System.
We now present our results for a five dimensional system following Definition 4 and using Lemma 2. We were able to find a degree five polynomial barrier certificate. Two examples of the projection of the certificate over two states of the system are shown in Figure 2. Table II displays the true dynamics used for simulation as well as the relevant vectors and matrices used for the SOS optimization.
Table I: Relevant Information for the 2D System
@@
Table II: Relevant Information for the 5D System

### v2

@@ 5. Numerical simulations
to incorporate the semi-algebraic set expressions for $w,A$ and $B$ for the optimization process based on condition (3). Figure 1 shows the resulting sublevel set of the barrier certificate for a two dimensional system along with the designated initial and unsafe states. Table 1 displays the true dynamics used for simulation (data collection) as well as the relevant vectors and matrices used for the SOS optimization.
Figure 1. Barrier Certificate for a 2D System.
Figure 2. Barrier Certificate Projections for a 5D System.
We now present our results for a five dimensional system following Definition 4 and using Lemma 2. We were able to find a degree five polynomial barrier certificate. Two examples of the projection of the certificate over two states of the system are shown in Figure 2. Table 2 displays the true dynamics used for simulation as well as the relevant vectors and matrices used for the SOS optimization.
Table 1. Relevant Information for the 2D System
@@
Table 2. Relevant Information for the 5D System